# From Majority to Minority: A Diffusion-based Augmentation for Underrepresented Groups in Skin Lesion Analysis

### v1

@@ Abstract
AI-based diagnoses have demonstrated dermatologist-level performance in classifying skin cancer. However, such systems are prone to under-performing when tested on data from minority groups that lack sufficient representation in the training sets. Although data collection and annotation offer the best means for promoting minority groups, these processes are costly and time-consuming. Prior works have suggested that data from majority groups may serve as a valuable information source to supplement the training of diagnosis tools for minority groups. In this work, we propose an effective diffusion-based augmentation framework that maximizes the use of rich information from majority groups to benefit minority groups. Using groups with different skin types as a case study, our results show that the proposed framework can generate synthetic images that improve diagnostic results for the minority groups, even when there is little or no reference data from these target groups. The practical value of our work is evident in medical imaging analysis, where under-diagnosis persists as a problem for certain groups due to insufficient representation.
@@ 1 Introduction
Current AI-assisted diagnostic systems demonstrate expert-level capability in classifying skin cancers, often identified visually. [6, 16, 2]. These systems can potentially contribute to teledermatology as diagnostic and decision-support tools, enhancing diagnostic accessibility in rural areas [3]. However, despite such success, recent studies have highlighted their susceptibility to under-diagnosing minority groups, such as those with underrepresented skin types, hindering their ability to generalize across different demographic groups [4, 11]. Although the majority group contains rich lesion information, directly training models for cross-color classification using this data is challenging due to the domain gap caused by varying skin types [24]. Prior research has suggested using synthetic images generated from majority groups to supplement the training for minority groups [19].
@@
Augmenting skin condition data with synthetic images has been explored, owing to its potential to address common challenges for skin lesion analysis, such as data privacy, imbalance, and scarcity. Notably, Generative Adversarial Networks (GANs) [10] and Diffusion Models (DMs) [5] have emerged as leading techniques for generating high-quality skin lesion images. While GANs have successfully produced photorealistic synthetic images, their generation is uncontrollable [9, 18]. On the other hand, DMs pre-trained on extensive web data have enabled higher controllability over image generation through the guidance of textual prompts, allowing for the creation of diverse and high-fidelity images of target skin conditions and types.
@@
Existing studies have used diffusion models to augment skin disease datasets and improve diagnosis for minority skin types using two public datasets: Diverse Dermatology Images (DDI) [4] and Fitzpatrick17k [11]. Each image in these datasets is annotated with skin type labels based on the Fitzpatrick scoring system [7]. In their work, [21] generated multiple synthetic images for each real image using Stable Diffusion [20] and then trained the classifier on a dataset including real and synthetic data. They find that diffusion models can enhance classifier performance across skin types in binary malignancy classification on the DDI dataset, though the number of real images remains a key driver in performance. Additionally, [22] sampled a small number of seed images with skin type at the end of the Fitzpatrick spectrum (FST I-II and FST V-VI) and carefully cropped the disease pathology, before generating synthetic data from the seed data with OpenAI DALL·E 2’s inpainting feature. They conducted class-wise data augmentation by incorporating synthetic images of the target condition and minority skin type into the real training set. Other related studies have focused on internal datasets [1, 15].
@@
Despite these advancements, the potential to leverage diffusion models’ knowledge about skin variation and the rich lesion information from majority groups to benefit minority groups remains underexplored. In this work, we propose a novel diffusion-based augmentation framework capable of learning skin lesion concepts from majority groups and generating images to improve classification performance for minority groups. Unlike current works that assume the existence of data from minority groups, we hypothesize that the information gained from majority groups and the diffusion model’s knowledge is sufficient to generate useful synthetic data. We test our hypothesis in a challenging multi-condition task. The framework is illustrated in Fig. 1. We conduct our experiments on the Fitzpatrick17k dataset, which includes lesions that are less familiar to diffusion models than common skin cancer. This dataset has a skewed skin type distribution, with light skin types (FST I-II) being significantly more common than dark skin types (FST V-VI), thus forming majority and minority groups. Our investigation focuses on images from both groups and is structured around three scenarios with increasing difficulty: (i) the training source includes some data from both groups; (ii) there is limited data from the minority group in the training source; and (iii) the training source lacks data from the minority group. Our findings reveal:
@@
  • Our proposed method effectively leverages lesion information from the majority group to generate synthetic images that can improve classification results for the minority group across all settings, even without reference data from the minority group.
@@
  • Using synthetic images generated by our method to train classifiers consistently outperforms training with only real images across various architectures. Further improvement is observed when combining real and synthetic data.
@@
Figure 1: Outline of the proposed augmentation framework. The framework pairs each training image with a textual prompt describing the condition as an input to train a latent diffusion model. Embeddings associated with new lesion concepts are found through Textual Inversion. Compact matrices $A$ and $B$ are optimized with LoRA to facilitate training with the new embeddings. During inference, the trained model produces synthetic images from the training data that mainly feature the majority groups via image-to-image generation, thus conditioned on visual cues of lesions from images and textual prompts describing the target condition and group attributes.
@@ 2 Methods
2.1 Latent Diffusion Models
@@
2.2 Concept Discovery via Textual Inversion
@@
Our proposed framework leverages Textual Inversion [8] to capture a unique embedding that accurately represents the targeted skin lesion concept from training data. Skin lesion images paired with a string containing a placeholder word (e.g., ‘An image of $\{S_{*}\}$’) are used to guide the learning of new lesion embedding for the generative model. In particular, The optimal embedding $v_{*}$ for the lesion concept $S_{*}$ is derived by minimizing the reconstruction loss
$v_{*}=\mathrm{argmin}_{v}\mathbb{E}_{z\sim\mathcal{E}(x),y,\epsilon\sim\mathcal{N}(0,1),t}\Big{[}\|\epsilon-\epsilon_{\theta}(z_{t},t,c_{\theta}(y))\|_{2}^{2}\Big{]}\,.$
This objective encourages the embedded representation to encapsulate new concepts specialized to skin disease datasets by re-using the same training scheme as the original LDM model while keeping both $c_{\theta}$ and $\epsilon_{\theta}$ fixed.
2.3 Fine-grained Detail Enhancement with LoRA
@@
To enhance efficiency in fine-tuning LDM, we employ Low-Rank Adaptation (LoRA) [14] in our framework, with the discovered tokens after textual inversion. This fine-tuning strategy freezes the pre-trained model weights and introduces two compact matrices $A$ and $B$, where $A\in\mathbb{R}^{n\times r},B\in\mathbb{R}^{r\times n}$. The adaptation matrices $AB$ are integrated into the attention layers to capture the fine visual details of the skin lesion concept with target embedding $v_{*}$. The optimization is formulated as
$L:=\mathbb{E}_{z\sim\mathcal{E}(x),y,\epsilon\sim\mathcal{N}(0,1),t}\Big{[}\|\epsilon-\epsilon_{\theta_{AB}}(z_{t},t,c_{\theta_{AB}}(v_{*}))\|_{2}^{2}\Big{]}\,.$
This step further strengthens the model to generate synthetic images that reflect fine-grained features not initially present in the pre-trained model.
@@
Architecture Train Accuracy Precision Recall F1 Score
VGG-16 real 70.24 $\pm$ 0.12 72.49 $\pm$ 0.37 69.58 $\pm$ 0.30 70.48 $\pm$ 0.13
syn 75.00 $\pm$ 0.64 75.77 $\pm$ 0.39 73.17 $\pm$ 0.29 72.42 $\pm$ 0.61
real + syn 77.98 $\pm$ 0.40 81.51 $\pm$ 0.25 78.87 $\pm$ 0.23 77.45 $\pm$ 0.29
ResNet-18 real 68.45 $\pm$ 0,42 69.42 $\pm$ 0.57 69.05 $\pm$ 0.42 67.94 $\pm$ 0.35
syn 69.05 $\pm$ 0.36 69.57 $\pm$ 0.82 69.05 $\pm$ 0.42 68.02 $\pm$ 0.33
real + syn 71.36 $\pm$ 0.69 69.02 $\pm$ 0.48 68.45 $\pm$ 0.40 67.56 $\pm$ 0.59
ViT-B-16 real 70.38 $\pm$ 0.42 73.72 $\pm$ 0.69 70.82 $\pm$ 0.39 70.61 $\pm$ 0.59
syn 74.19 $\pm$ 0.37 77.89 $\pm$ 1.01 74.04 $\pm$ 0.85 73.58 $\pm$ 0.53
real + syn 78.65 $\pm$ 0.53 81.57 $\pm$ 0.47 79.17 $\pm$ 0.84 78.24 $\pm$ 0.64
@@
Table 2: This table presents the results when the training set includes some (291 non-seed) images from the minority group. The test set comprises 56 seed images of the minority group (FST V-VI), uniformly distributed across the 7 conditions. Here, “real” indicates that the classifier is trained solely on real images, while “syn” means that it is trained exclusively on synthetic images generated by our framework. Accordingly, “real+syn” means that the subsequent classifier is trained on a combination of both.
@@ 3 Experiments
We conduct our experiments using the Fitzpatrick17k dataset, where each image is annotated with a condition and a Fitzpatrick Skin Type (FST) label. In line with [22], our analysis narrows down to a subset of the Fitzpatrick17k dataset, encompassing 7 conditions (Table 1). These conditions were selected because they represent the largest sample sizes at the ends of the Fitzpatrick Skin Type (FST I-II or V-VI) spectrum. Unlike [22], our study excludes intermediate skin types (FST III-IV), to explore the efficacy of our diffusion-based augmentation in a more challenging and explainable way. For each condition, we randomly sample 8 images from the lightest (FST I-II) and darkest (FST V-VI) skin type groups, resulting in 56 seed images for each group.
@@
We independently examine three scenarios: (i) the training source includes some images of both dark and light skin types, and the target features uniformly distributed seed images (56 of one skin type) over the 7 conditions; (ii) there are many light-skinned images and a few dark-skinned seed images in the training set. In this setting, the target consists of 291 holdout images of dark skin type; and (iii) the source lacks dark-skinned images entirely and includes only light skin type. Likewise, the target comprises of 291 dark-skinned images. In all the settings, we generate 5 synthetic images for each real one in the source training set during inference, using the fine-tuned model, as illustrated by Fig. 1.
In setting (i), to ensure a sufficient number of examples for both majority and minority groups (lightest vs. darkest skin types), we designate the seed images as the target and utilize all remaining images as sources for generator and classifier training. Meanwhile, this setting also serves as the basis for hyperparameter tuning of the diffusion model, with the chosen hyperparameters being fixed for subsequent experiments. In setting (ii), we reversed the roles of the seed and non-seed images, creating a training set composed of 56 dark-skinned seed images with other light-skinned ones, and a test set of 291 dark-skinned images. In the final setting, we removed the seed images from the training set to evaluate whether our pipeline could still generate synthetic images that enhance classification performance for the target minority groups, even without any visual references from them. Note that the same set of images is used for training the diffusion model and classifier in all settings.
@@
Implementation Details In each setting, we sampled 5 groups of seed images and repeated the experiment 5 times. We used the Stable Diffusion 2.1 base [20] and the Diffusers library [17] for fine-tuning the diffusion model and generating synthetic images. We used pre-trained VGG-16 [23], ResNet-18 [12], and ViT-B-16 [25] as classifier backbones and trained each classifier for 30 epochs using the Adam optimizer with an initial learning rate of 1e-3. A weight-based sampler and StepLR scheduler were applied. Experiments were conducted on two NVIDIA GeForce RTX 3090s.
@@
Architecture Train Accuracy Precision Recall F1 Score
VGG-16 real 58.79 $\pm$ 0.10 58.90 $\pm$ 0.03 58.26 $\pm$ 0.05 56.98 $\pm$ 0.04
syn 62.86 $\pm$ 0.15 61.49 $\pm$ 0.15 63.32 $\pm$ 0.13 61.57 $\pm$ 0.24
real + syn 63.66 $\pm$ 0.11 62.80 $\pm$ 0.12 64.08 $\pm$ 0.08 62.72 $\pm$ 0.18
ResNet-18 real 50.31 $\pm$ 0.30 50.45 $\pm$ 0.44 51.27 $\pm$ 0.30 48.23 $\pm$ 0.32
syn 56.36 $\pm$ 0.16 56.58 $\pm$ 0.13 59.78 $\pm$ 0.08 55.58 $\pm$ 0.10
real + syn 61.33 $\pm$ 0.17 59.75 $\pm$ 0,12 62.52 $\pm$ 0.17 59.94 $\pm$ 0.17
ViT-B-16 real 62.03 $\pm$ 0.19 63.09 $\pm$ 0.13 62.02 $\pm$ 0.04 61.05 $\pm$ 0.04
syn 68.83 $\pm$ 0.19 70.07 $\pm$ 0.03 68.22 $\pm$ 0.25 68.34 $\pm$ 0.19
real + syn 71.20 $\pm$ 0.07 71.61 $\pm$ 0.19 71.53 $\pm$ 0.01 71.17 $\pm$ 0.13
@@
Table 3: Classification results when the training set contains few reference images from the minority group (56). The test set has 291 non-seed images of the minority group (FST V-VI).

### v2

@@ Abstract
AI-based diagnoses have demonstrated dermatologist-level performance in classifying skin cancer. However, such systems are prone to under-performing when tested on data from minority groups that lack sufficient representation in the training sets. Although data collection and annotation offer the best means for promoting minority groups, these processes are costly and time-consuming. Prior works have suggested that data from majority groups may serve as a valuable information source to supplement the training of diagnostic tools for minority groups. In this work, we propose an effective diffusion-based augmentation framework that maximizes the use of rich information from majority groups to benefit minority groups. Using groups with different skin types as a case study, our results show that the proposed framework can generate synthetic images that improve diagnostic results for the minority groups, even when there is little or no reference data from these target groups. The practical value of our work is evident in medical imaging analysis, where under-diagnosis persists as a problem for certain groups due to insufficient representation. Our implementation detail is available at https://github.com/janet-sw/skin-diff.
@@ 1 Introduction
AI-assisted diagnostic systems demonstrate expert-level capability in classifying skin cancers, often identified visually [6, 16, 2]. These systems can potentially contribute to teledermatology as diagnostic and decision-support tools, enhancing diagnostic accessibility in rural areas [3]. However, despite such success, recent studies have highlighted their susceptibility to under-diagnosing minority groups, such as those with underrepresented skin types, hindering their ability to generalize across different demographic groups [4, 11]. Although the majority group contains rich lesion information, directly training models for cross-color classification using this data is challenging due to the domain gap caused by varying skin types [24]. Prior research has suggested using synthetic images generated from majority groups to supplement the training of AI for minority groups [19].
@@
Augmenting skin condition data with synthetic images has been explored, owing to its potential to address common challenges for skin lesion analysis, such as data privacy, imbalance, and scarcity. Notably, Generative Adversarial Networks (GANs) [10] and Diffusion Models (DMs) [5] have emerged as leading techniques for generating high-quality skin lesion images. While GANs have successfully produced photorealistic synthetic images, their generation is uncontrollable [9, 18]. On the other hand, DMs pre-trained on extensive web data have enabled higher controllability over image generation through the guidance of textual prompts, allowing for the creation of diverse and high-fidelity images of target skin conditions and skin types.
@@
Existing studies have tried diffusion models to augment minority skin types using two public datasets: Diverse Dermatology Images (DDI) [4] and Fitzpatrick17k [11]. Each image in these datasets is annotated with skin type labels based on the Fitzpatrick scoring system [7]. In their work, [21] generated multiple synthetic images for each real image using Stable Diffusion [20] and then trained the classifier on a dataset including real and synthetic data. They found that diffusion models can enhance diagnosis accuracy across skin types in binary malignancy classification on the DDI dataset, though the number of real images is the key driver in performance. Additionally, [22] sampled a small number of seed images with skin types at the ends of the Fitzpatrick spectrum (FST I-II and FST V-VI) and carefully cropped the disease pathology in them, before generating synthetic data from the seeds using OpenAI DALL·E 2’s inpainting feature. They conducted class-wise data augmentation by incorporating synthetic images of the target condition and minority skin type into the real training set. Other related studies have focused on internal datasets [1, 15].
@@
Despite these advancements, the potential to leverage diffusion models’ knowledge about skin variation and the rich lesion information from majority groups to benefit minority groups remains underexplored. In this work, we propose a novel diffusion-based augmentation framework capable of learning skin lesion concepts from majority groups and generating images to improve classification performance for minority groups. Unlike current works that assume the existence of data from minority groups, we hypothesize that the information gained from majority groups and the diffusion model’s pre-trained knowledge is sufficient to generate useful synthetic data. We test our hypothesis in a challenging multi-condition classification task. The framework is illustrated in Fig. 1. We conduct our experiments on the Fitzpatrick17k dataset, which includes lesions that are less familiar to diffusion models than common skin cancer. This dataset has a skewed skin type distribution, with light skin types (FST I-II) being significantly more than dark skin types (FST V-VI), thus forming majority and minority groups. Our investigation focuses on images from both groups and is structured around three scenarios with increasing difficulty: (i) the training source includes some data from both groups; (ii) there is limited data from the minority group in the training source; and (iii) the training source lacks data from the minority group. Through extensive experiments and analysis, we found that:
@@
  • Our proposed method effectively leverages lesion information from the majority group to generate synthetic images that can improve classification for the minority group across all settings, even without reference data from the minority group.
@@
  • Using synthetic images generated by our method to train classifiers consistently outperforms training with real images across various architectures. Further improvement is observed when combining real and synthetic data.
@@
Figure 1: Overview of the proposed augmentation framework. The framework pairs each training image with a textual prompt describing the condition as an input to train a latent diffusion model. Embeddings associated with new lesion concepts are found through Textual Inversion. Compact matrices $A$ and $B$ are optimized via LoRA to facilitate training with the new embeddings. During inference, the trained model produces synthetic images from the training set that mainly features the majority groups via image-to-image generation, thus conditioned on visual cues of lesions from images and textual prompts describing the target condition and group attributes.
@@ 2 Methods
In this section, we will introduce key techniques that have been adapted for skin disease datasets in our proposed augmentation framework.
@@ 2.0.1 Latent Diffusion Models
2.0.2 Concept Discovery via Textual Inversion
@@
Our proposed framework leverages Textual Inversion [8] to capture a unique embedding that accurately represents the targeted skin lesion concept from training data. Skin lesion images paired with a string containing a placeholder word (e.g., ‘An image of $\{S_{*}\}$’) are used to guide the learning of a new lesion embedding for the generative model. In particular, the optimal embedding $v_{*}$ that encapsulates the lesion concept $S_{*}$ is derived by minimizing the reconstruction loss,
$v_{*}=\mathrm{argmin}_{v}\mathbb{E}_{z\sim\mathcal{E}(x),y,\epsilon\sim\mathcal{N}(0,1),t}\Big{[}\|\epsilon-\epsilon_{\theta}(z_{t},t,c_{\theta}(y,S_{*}))\|_{2}^{2}\Big{]}\,,$
where the same training scheme as the original LDM model is used, with $c_{\theta}$ and $\epsilon_{\theta}$ fixed.
2.0.3 Fine-grained Detail Enhancement with LoRA
@@
To enhance efficiency in fine-tuning LDM, we employ Low-Rank Adaptation (LoRA) [14] in our framework, with the discovered tokens after textual inversion. This fine-tuning strategy freezes the pre-trained model weights and introduces two compact matrices $A$ and $B$, where $A\in\mathbb{R}^{n\times r},B\in\mathbb{R}^{r\times n}$. The adaptation matrices $AB$ are integrated into the attention layers to capture fine visual details of the skin lesion that were not initially present in the pre-trained model, with target embedding $v_{*}$. The optimization is formulated as
$L:=\mathbb{E}_{z\sim\mathcal{E}(x),y,\epsilon\sim\mathcal{N}(0,1),t}\Big{[}\|\epsilon-\epsilon_{\theta_{AB}}(z_{t},t,c_{\theta_{AB}}(y,v_{*}))\|_{2}^{2}\Big{]}\,.$
@@
     Architecture Train Type Train Size Accuracy Precision Recall F1 Score
VGG-16 real 1519 70.24 $\pm$ 0.12 72.49 $\pm$ 0.37 69.58 $\pm$ 0.30 70.48 $\pm$ 0.13
syn 1519 * 5 75.00 $\pm$ 0.64 75.77 $\pm$ 0.39 73.17 $\pm$ 0.29 72.42 $\pm$ 0.61
real + syn 1519 * 6 77.98 $\pm$ 0.40 81.51 $\pm$ 0.25 78.87 $\pm$ 0.23 77.45 $\pm$ 0.29
ResNet-18 real 1519 68.45 $\pm$ 0,42 69.42 $\pm$ 0.57 69.05 $\pm$ 0.42 67.94 $\pm$ 0.35
syn 1519 * 5 69.05 $\pm$ 0.36 69.57 $\pm$ 0.82 69.05 $\pm$ 0.42 68.02 $\pm$ 0.33
real + syn 1519 * 6 71.36 $\pm$ 0.69 69.02 $\pm$ 0.48 68.45 $\pm$ 0.40 67.56 $\pm$ 0.59
ViT-B-16 real 1519 70.38 $\pm$ 0.42 73.72 $\pm$ 0.69 70.82 $\pm$ 0.39 70.61 $\pm$ 0.59
syn 1519 * 5 74.19 $\pm$ 0.37 77.89 $\pm$ 1.01 74.04 $\pm$ 0.85 73.58 $\pm$ 0.53
real + syn 1519 * 6 78.65 $\pm$ 0.53 81.57 $\pm$ 0.47 79.17 $\pm$ 0.84 78.24 $\pm$ 0.64
@@
Table 2: This table presents the results when the training set includes non-flexible images from the minority group (291 of FST V-VI) and the majority group (1228 FST I-II). The test set is a flexible subset of the minority group (56 of FST V-VI), uniformly distributed across the 7 conditions. Here, “real" indicates that the classifier is trained solely on real images, while “syn" means that it is trained exclusively on synthetic images generated by our framework. Accordingly, “real+syn" means the subsequent classifier is trained on a combination of both.
@@ 3 Experiments
We conduct our experiments using the Fitzpatrick17k dataset, where each image is annotated with a condition and a Fitzpatrick Skin Type (FST) label. In line with [22], our analysis narrows down to a subset of the Fitzpatrick17k dataset, encompassing 7 conditions (Table 1). These conditions were selected because they represent the largest sample sizes at the ends of the Fitzpatrick Skin Type (FST I-II or V-VI) spectrum. Unlike [22], our study excludes intermediate skin types (FST III-IV), to explore the efficacy of our diffusion-based augmentation in a more challenging and explainable way. We randomly sample 8 images for each condition from the lightest (FST I-II) and darkest (FST V-VI) skin type groups, resulting in a flexible subset of 56 images for each group.
@@
We examine three scenarios: (i) the training set includes images of dark and light skin types, and the test set features the uniformly distributed flexible subset across the 7 conditions; (ii) the training set predominantly includes light-skinned images and a few dark-skinned images, while the test set consists of dark-skin data; (iii) the training set lacks dark-skinned images entirely, while the test set comprises dark-skinned images. In all scenarios, we generate 5 synthetic images for each real one in the training set during inference, using the fine-tuned model, as illustrated in Fig. 1. In scenario (i), to ensure a sufficient number of examples for both majority and minority groups in the training set, we designate the flexible subset of dark skin as the test set and use remaining non-flexible images for generator and classifier training. This setting also serves as the basis for hyperparameter tuning of the diffusion model, with the selected hyperparameters being fixed for subsequent experiments.
@@
Implementation Details In each setting, we randomly sampled 5 flexible subsets and repeated the experiment 5 times. We used the Stable Diffusion 2.1 base [20] and the Diffusers library [17] for fine-tuning the diffusion model and generating synthetic images. For classifier backbones, we utilized pre-trained VGG-16 [23], ResNet-18 [12], and ViT-B-16 [25] and trained each classifier using the Adam optimizer with an initial learning rate of 1e-3 and transformations as in [11]. A weight-based sampler and StepLR scheduler were applied. All experiments were conducted on two NVIDIA GeForce RTX 3090s.
@@
     Architecture Train Type Train Size Accuracy Precision Recall F1 Score
VGG-16 real 1284 58.79 $\pm$ 0.10 58.90 $\pm$ 0.03 58.26 $\pm$ 0.05 56.98 $\pm$ 0.04
syn 1284 * 5 62.86 $\pm$ 0.15 61.49 $\pm$ 0.15 63.32 $\pm$ 0.13 61.57 $\pm$ 0.24
real + syn 1284 * 6 63.66 $\pm$ 0.11 62.80 $\pm$ 0.12 64.08 $\pm$ 0.08 62.72 $\pm$ 0.18
ResNet-18 real 1284 50.31 $\pm$ 0.30 50.45 $\pm$ 0.44 51.27 $\pm$ 0.30 48.23 $\pm$ 0.32
syn 1284 * 5 56.36 $\pm$ 0.16 56.58 $\pm$ 0.13 59.78 $\pm$ 0.08 55.58 $\pm$ 0.10
real + syn 1284 * 6 61.33 $\pm$ 0.17 59.75 $\pm$ 0,12 62.52 $\pm$ 0.17 59.94 $\pm$ 0.17
ViT-B-16 real 1284 62.03 $\pm$ 0.19 63.09 $\pm$ 0.13 62.02 $\pm$ 0.04 61.05 $\pm$ 0.04
syn 1284 * 5 68.83 $\pm$ 0.19 70.07 $\pm$ 0.03 68.22 $\pm$ 0.25 68.34 $\pm$ 0.19
real + syn 1284 * 6 71.20 $\pm$ 0.07 71.61 $\pm$ 0.19 71.53 $\pm$ 0.01 71.17 $\pm$ 0.13
@@
Table 3: Classification results when the training set contains a few reference images from the dark-skinned flexible subset (56 of FST V-VI) and non-flexible light-skinned images (1228 of FST I-II). The test set includes all other dark-skinned images (291 of FST V-VI) outside the flexible subset.